OpenStudy (anonymous):

write an equation of the line containing the given point and parallel to the given line. Express you answer in the form of y=mx +b. (-3, 9); 2x= 3y + 5.

OpenStudy (cathyangs):

if it's parallel to the given line, they have the same slope. You can use point-slope form and then some algebra to get to slope-intercept.

OpenStudy (mathmate):

2(x-(-3))=3(y-9) 2x+6=3y-27 3y=2x+33 y=(2/3)x+11